OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4)

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

y-y1/y2-y1=x-x1/x2-x1 put points in equtaion and solve

OpenStudy (anonymous):

i have a tricjk u can see both the points have same y co rdinate tus the line must be y=4

OpenStudy (anonymous):

4-4/4-4=5-2/5-2 @03225186213 what do i do now?

OpenStudy (anonymous):

0/0=3/3? @03225186213

OpenStudy (anonymous):

thats why i told u to do my way actually 0/0 is undefined and undefined slope means it is y axis

OpenStudy (anonymous):

@thinker1 but was i right?

OpenStudy (anonymous):

y and x are not given y-4/4-4=x-2/5-2 (y-4)(3)=0 3y-12=0

OpenStudy (anonymous):

it is horizomtal line as slope =0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

no see 03225186213 has solved it for u but i am saying that u could have directly taken the line to be y=4 as u can see same ordinates of 2 point this saves ur time

OpenStudy (anonymous):

well ok thanks :)

OpenStudy (anonymous):

u r welcome